a quadratic function has an equation that can be written in the form f(x)=a(x-r)(x-s). the graph of the function has x-intercepts at (1,0) ans (3,0) and passes through the point (-1, 16). write the equation of the function in factored and then in standard form
Answer by Alan3354(69443)

a quadratic function has an equation that can be written in the form f(x)=a(x-r)(x-s). the graph of the function has x-intercepts at (1,0) and (3,0) and passes through the point (-1,16). write the equation of the function in factored and then in standard form
--------------
f(x) = a(x-1)*(x-3) = ax^2 - 4ax + 3a
---
f(-1) = a + 4a + 3a = 8a = 16
a = 2
==============
f(x) = 2x^2 - 8x + 6